a partyicle covers 10m in first 5s and 10m in next 3s. assuming constant acceleration, find intail speed, acceleration and distance covered in next 2s.

Grade:9

1 Answers

PRAJVAL98
9 years ago
use s=ut + (at^2)/2
          u get 2 eqs with 2 variables u & a
         for the 1st eq use t=5s & 2nd t=8s
         u get the following eq
10=5u+12.5a  and
 20(i.e s1+ s2)=8u+32a         here t=8s
u get a=1/3 & u=1.667
 
 
let time for next 2s be t3=10s
s= 10*1.67+100*0.33/2
s=16.7+16.5
s=32.5